OpenStudy (anonymous):

how many solutions does the eqn |3x+7|+2=0 have ? please explain .

OpenStudy (anonymous):

x= -3,-5/3

OpenStudy (anonymous):

please explain

OpenStudy (nubeer):

3x+7 = -2 & -(3x+7) = -2 just solve these 2 equation u will get 2 values of x

OpenStudy (anonymous):

so it will always be two ?

OpenStudy (anonymous):

2 solutions that is

OpenStudy (nubeer):

when this mod sign is applied always 2 l l